OpenStudy (anonymous):

What fraction is the same as 62/8?

OpenStudy (anonymous):

7 and 6/8

OpenStudy (anonymous):

or for just a fraction 31/4

OpenStudy (anonymous):

how you find that out ?

OpenStudy (anonymous):

You just divide the fraction

OpenStudy (anonymous):

You take a common factors of 62 and 8, and divide both by it. In this case, they have a common factor of 2. This process is called "simplifying" the fraction. 62/2 = 31 8/2 = 4 So, 62/8 = 31/4